For the purple section shown below, determine the orientation of the principal centroidal axes in degrees and the principal centroidal moments of inertia in mm4. The thickness of each rectangle is 10 mm. Use Mohr's Circle.

650 mm 630 mm 660 mm 640 mm 650 mm 660 mm mm4 min mm4 Imax =
Answer #1

Sotu 2 2. 2 2 2 2 20 +10. = 25 mm 2.87. I g4 mm 2 1 2 1 2.2. 12 8,172x lo mm2 2 8.428 x 108 mm 21:44x108 +2556 4 x lo8 + g , y 28 x108 = 2 31.93x lo mm = 22- 2 Tr tan 2.gf.= = 8.172x108